Humans who are born homozygous for the recessive sickle cell allele die of sickle cell anemia, while those who are heterozygous are resistant to malaria (see chapter 4 for more information on this balanced polymorphism).  4% of the population of the Congo are homozygous recessive for the sickle cell allele.  What percentage of the population is resistant to malaria?


Red hair is a result of a recessive mutation at the melanocortin 1 receptor (MC1R).  As you learned, about 13% of the Scottish population is red haired. If the population mates randomly, what is the frequency of heterozygotes?


In the Ireland, 10% of the population is red haired.  Assuming mating is random, what is the frequency of heterozygotes in Ireland?
Answer #1

Answer:

Given-

Frequency of homozygous recessive (q^2)=4/100=0.04

So q=√0.04=0.2

According to Hardy Weinberg equation p+q=1

So p=1-0.2=0.8

To know the heterozygous population (resistant) 2pq= 2×0.8×0.2= 0.32

Consequently 32 percent of people are resistant to malaria.

**** Please post the remaining questions separately i couldn't answer all due to time consent. Give a thumbs up if you like the answer. Thank you